OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally144cubic meters and is then resized using a scale factor of1/2on each side length?

OpenStudy (anonymous):

help me please

OpenStudy (unklerhaukus):

your question has some stuff missing

OpenStudy (anonymous):

fixed

OpenStudy (unklerhaukus):

The Volume of a rectangular prims is \[V=l\times w\times h\]

OpenStudy (unklerhaukus):

if each of l, w, and h are halved \[\frac l2\times\frac w2 \times\frac h2=\frac 1 8V\]

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

what is the new volume is cubic meters?

OpenStudy (anonymous):

\[28m^3\]

OpenStudy (anonymous):

is it right

OpenStudy (unklerhaukus):

\[\frac{144m^3}{8}\]

OpenStudy (anonymous):

i still do no quite understand could someone help me some more

OpenStudy (anonymous):

I would really appreciate it

OpenStudy (unklerhaukus):

144/8=18